When graphing a line that has slope 2/3 and passes through the point (1, -4, ehich of the following is a second point on the line?
A) (3,-1) B) (4-2) C) (-1,1) D) (3,8) E) (2,3)

Answer by checkley71(8403)

Y=mX+b
-4=2/3*1+b
-4=2/3+b
b=-4-2/3
b=-14/3 THUS WE HAVE A Y INTERCEPT OF -14/3
Y=2/3X-14/3
-2=2/3*4-14/3
-2=8/3-14/3
-2=-6/3
-2=-2 PROOF THAT B)4,-2 IS A POINT ON THIS LINE.